Operations Report – April 22nd

Crew 297 Operations Report 22-04-2024

SOL: 8

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve), Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. Suit 11’s battery was replaced yesterday, but on today’s EVA Suit 11 was believed to exhibit reduced circulation (albeit better than before the battery was replaced). After inspection the cause remains unknown. 3 missing zip ties for the tarps in the tunnels were replaced.


Opportunity rover used:

Hours: 183.4

Beginning charge: 100%

Ending charge: 46%

Currently charging: No

General notes on rovers: NA

Summary of Hab operations:

WATER USE: 25.4 gallons

Water (static tank): 233.4 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: Looking up recipes, downloading entertainment, catching up on Earthly news

Summary of suits and radios:

Radios 1,5, and 6 were used.

EVA Suits 4 and 7 were used.

Summary of GreenHab operations:

WATER USE: 4 gallons

Heater: Yes

Supplemental light: N/A

Harvest: 2g of cherry tomatoes

Summary of ScienceDome operations: Setup for Pawel’s EVA tomorrow took place in the ScienceDome now that the noxious smell is gone.

Summary of RAM operations: (Tools used, work done): 3 missing zip-ties for the tarps on the tunnels were replaced using the supplies from the RAM.

Summary of any observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site